# Individualizing Glioma Radiotherapy Planning by Optimization of a Data and Physics-Informed Discrete Loss

### v3

@@ 2.2 Radiotherapy Planning and Recurrence coverage
Our main analysis encompasses two distinct definitions of tumor recurrence: a broad definition including edema, enhancing core, and necrotic core, and a more specific definition aligned with current RANO guidelines focusing only on the enhancing core. Figures 5 and 6 present the mean and standard deviation of Recurrence Coverage for these definitions. Furthermore, we conduct a patient-specific comparison with the Standard Plan, classifying a model as ’Better’ if it provides higher coverage for an individual patient and ’Worse’ if it falls short of the Standard Plan’s coverage. In instances of equal coverage (e.g., both achieving 100%), we label the outcome as ’Equal’. The comparison outcomes, including average results, are depicted in Figures 5c and 6c.
Significant disparities are evident between the outcomes from any segmentation recurrence and those from enhancing core recurrence only. First, let’s examine the results from radiotherapy targeting any segmentation recurrence. GliODIL demonstrates superior performance over uniform margin plans in most instances. Specifically, for the 58 patients with available FET-PET scans, the average Recurrence Coverage improved from 71.85% to 76.09%. For the 152 patients without FET-PET scans or where the scans were not incorporated into the model, the coverage increased from 64.56% to 70.31%. These improvements translate to a 35% and 38% difference in patient outcomes favoring GliODIL over the standard approach for the respective datasets.
Figure 5c reveals significant improvements, with the percentage of patients benefiting from the treatment increasing from 21% to 35% when comparing GliODIL, enhanced with FET-PET imaging, to the Standard Radiotherapy Plan with a 1.5cm margin. This emphasizes the value of integrating FET-PET modality into flexible models capable of utilizing multi-modal data. Conversely, incorporating this modality into models rigidly conforming to a basic predefined PDE family yields statistically insignificant changes in performance.
@@
In the analysis of enhancing core recurrence in follow-up scans, the hierarchy observed in the any segmentation recurrence study largely remains, with the exception that the Standard Plan demonstrates significantly improved performance, outshining all models strictly based on PDE approaches. Instances of equal outcomes, as depicted in Figures 6c,d, become more prevalent, indicating that both compared methods often achieve 100% coverage.
GliODIL stands out as the only model surpassing the Standard Plan, especially notable in a cohort of patients where FET-PET imaging is utilized. This leads to a marked increase in coverage improvement, from benefiting 5% of patients to 19%. The conclusion drawn is that it’s considerably more challenging for models to surpass the uniform margin around the pre-operative visible tumor. This difficulty arises because recurrences to the enhancing core typically occur closer to the resection cavity, rendering the migration paths, which these models leverage, less critical in this context compared to any segmentation recurrence. Any segmentation recurrences often happen further away, where anatomy and topological barriers, not considered by the uniform margin, play a significant role. GliODIL, particularly when incorporating FET-PET, can utilize additional information while still adhering closely to pre-operative data, thereby outperforming the standard margin approach.
Figure 5: Real patient data. Recurrence of edema, enhancing/necrotic core coverage overview. a,b Recurrence coverage of selected volume radiotherapy plans. All radiotherapy plans have the same total volume. Output tumor cell distribution thresholds found through a grid search to match the Standard Plan volumes. c,d Average Recurrence Coverage and direct patient-by-patient comparisons to the Standard Plan.
Figure 6: Real patient data. Recurrence of the enhancing core overview. a,b Recurrence coverage of selected volume radiotherapy plans. All radiotherapy plans have the same total volume. Output tumor cell distribution thresholds found through a grid search to match the Standard Plan volumes. c,d Average Recurrence Coverage and direct patient-by-patient comparisons to the Standard Plan.
@@
Ending this section, we note that the $\text{PDE}_{\text{GliODIL}}$ model, through the comparative analysis detailed in Section 2.1, aligns with the $\text{PDE}_{\text{CMA-ES}}$ in capturing pre-operative data yet surpasses it in recurrence prediction accuracy. This suggests the parameter estimations by $\text{PDE}_{\text{GliODIL}}$ more accurately reflect patient-specific conditions. Such improvements imply the data-driven component of GliODIL benefits the fine-tuning of PDE parameters which can be used for both diagnostics and forcasting.
@@ 3 Discussion
Leveraging its ability to synthesize information from diverse modalities, GliODIL significantly benefits from the integration of FET-PET imaging, which notably enhances its predictive accuracy beyond the capabilities of the Standard Radiotherapy Plan. This plan typically employs uniform safety margins around the visible tumor regions identified on MRI scans, aiming to encapsulate both the visible tumor and potential microscopic tumor extensions not visible on imaging. This approach, while clinically practiced, lacks the specificity required for optimal therapeutic outcomes. The improvements GliODIL brings to the table, as evidenced by the patient outcomes depicted in Figures 5 and 6, underscore the framework’s superior capability to utilize complex imaging data to refine treatment strategies. Such performance enhancements are not achieved by models strictly adhering to PDE solutions, which struggle to capture the full extent of tumor recurrence, even with the addition of extra imaging inputs.
GliODIL’s adaptive modeling approach provides a stark contrast to the Standard Radiotherapy Plan’s reliance on uniform margins, offering a much-needed precision in targeting the tumor. Moreover, GliODIL is adept at incorporating not just FET-PET imaging data but also embracing other diagnostic modalities such as MR diffusion imaging. This particular MR technique provides valuable insights into tissue cellularity, potentially further enriching the model’s understanding of the tumor environment.
Our results thus highlight both the relevant contribution of advanced, biological imaging techniques to inform about the underlying tumor biology and the ability of our GliODIL framework to flexibly incorporate such additional information to improve inverse problem-solving validated by the clinically important tumor recurrence prediction task. Our current study is concentrated on immediate post-operative treatment, necessitating reliance on a single imaging time point. In scenarios where time-series data are accessible, the exciting opportunity arises to employ multi-mixture PDE models, which, despite their challenges with just one time-point due to ill-posedness, become viable with the availability of sequential imaging data.
GliODIL can be extended to consider more complex physics models, including those accounting for mass effects through dynamical tissues and tackling the challenge of unknown initial healthy brain anatomy, all while maintaining the framework’s flexibility and computational efficiency. Addressing uncertainties in parameter inference remains a pivotal focus, potentially achievable by integrating variational inference techniques into the framework.
@@ 4.2 Optimizing a Discrete Loss (ODIL)
The multi-grid ODIL technique, introduced in the paper [26], builds upon the original ODIL methodology by incorporating a multigrid decomposition scheme to fasten the convergence process. This technique is particularly adept at leveraging the multi-scale attributes of the forward problem. It decomposes the problem into various scale bands, each characterized by different levels of detail. Formally, for a uniform grid with dimensions $N_{1}=N$, a hierarchical sequence of coarser grids is introduced with sizes defined as $N_{i}=N/2^{i-1}$ for $i=1,\ldots,L$.
@@
As depicted in Figure 1c, which illustrates the multigrid domain, a tumor growth is simulated over an ensemble of Cartesian 4D grids in both time and space, with each grid level being coarser than the preceding one. This hierarchical decomposition allows the optimization algorithm to initially concentrate on the coarse-scale features, incrementally incorporating finer-scale details as the process evolves. Such an approach not only enables a more comprehensive exploration of the parameter space but also sidesteps the pitfalls of local minima and expedites convergence.
We focus the computational grids on the tumor region. This leads to an average resolution of $71\times 68\times 55$ for the area of interest in our images, significantly reduced from the original $240\times 240\times 155$. For parameter inference within the GliODIL framework, we employed a $48^{3}$ spacial resolution and 192 degrees in the temporal resolution. Our tests indicated that this resolution is sufficient, revealing no significant differences compared to inference using the native resolution. For a forward PDE run with inferred parameters by GliODIL, which we call $\text{PDE}_{\text{GliODIL}}$, we use the native resolution.
@@ 4.3 Imaging Model
$L_{\text{DATA}}=\lambda_{\text{CORE}}L_{\text{CORE}}(u,\theta_{\text{up}})+\lambda_{\text{EDEMA}}L_{\text{EDEMA}}(u,\theta_{\text{down}},\theta_{\text{up}})+\lambda_{\text{PET}}L_{\text{PET}}(u,s,A_{\text{PET}})$ (10)
@@ 4.4 Final Loss Function
The final loss function measures discrepancy between proposed tumor cell concentrations $u(x,y,z,t)$ and our objective. It is a composite term that comprises contributions from multiple sources: the PDE constraint, data fitting, and additional regularization term. Specifically, the PDE loss, denoted by $L_{\text{PDE}}$ (introduced in 4.2), imposes the discretized PDE equation constraint. Initial condition loss term $L_{\text{IC}}$ in the overall loss function serves to enforce that the tumor at $t=0$ originates from a Gaussian origin (introduced in 4.1). The data loss (introduced in 4.3), denoted by $L_{\text{DATA}}=\lambda_{\text{CORE}}L_{\text{CORE}}+\lambda_{\text{EDEMA}}L_{\text{EDEMA}}+\lambda_{\text{PET}}L_{\text{PET}}$, accounts for matching tumor core and edema segmentations as well as fitting to PET metabolic data. Additional regularization term confines the inferred parameters within a plausible range $L_{\text{PARAMS}}$[29].
@@
where $\lambda_{*}$ are detailed on in Section 2.1 experiments.
@@ 4.5 Evaluation Metrics
Dice $\text{Dice Coefficient}=\frac{2\times|A\cap B|}{|A|+|B|}$ (18)
@@
The principal metric for evaluating our model’s effectiveness in radiotherapy planning is its accuracy in predicting tumor recurrence within the specified radiation volume after treatment. We identify the area of tumor recurrence using two definitions: a narrow definition focusing on the enhancing core observed in post-treatment patient data, and a broader definition that includes edema and the enhancing/necrotic core in the post-treatment follow-up MRI scans. We register the post-treatment patient anatomy to the pre-treatment in order to remove special shifts due to surgeries. The textbfRecurrence Coverage [%] metric refers to the percentage of segmented recurrences, that are encompassed within the radiation volume delineated by a given radiotherapy plan. Each radiotherapy plan maintains a uniform total volume, equivalent to that of the Standard Plan, which applies a consistent 1.5cm margin around the pre-operative segmented enhancing and necrotic regions. To facilitate a fair comparison, we adjust the Standard Plan’s total volume by excluding portions that extend beyond the patient’s anatomical boundaries before making any calculations.
4.6 Datasets
@@
We selected 152 adult patients from the glioma database at TUM University Hospital to validate our method in actual patient data. All patients were diagnosed with a WHO-CNS grade 4 IDH wild type glioblastoma, according to the 2021 WHO classification of brain tumors. Imaging data included a preoperative and postoperative MRI, as well as an MRI scan at first tumor recurrence following combined radio-chemotherapy according to the Stupp protocol. MR scans were performed on a 3T Philips MRI scanner (either Achieva or Ingenia; Philips Healthcare, Best, The Netherlands) and comprised 3D-T1w-MPRAGE images before and after administration of Gadolinium-based contrast agent, 3D-FLAIR images and 3D-T2w images (all 1mm isotropic voxel resolution). In addition, for 58 patients, preoperative FET-PET imaging was also available. FET-PET data were acquired on either a PET/MR (Biograph mMR, Siemens Healthcare GmbH, Erlangen, Germany) or a PET/CT (Biograph mCT; Siemens Healthcare, Knoxville, TN, USA), according to a standard clinical protocol. Patients were asked to fast for a minimum of 4 h before scanning. Emission scans were acquired at 30 to 40 min after intravenous injection of a target dose of 185$\pm$10 MBq [18F]-FET. Attenuation correction was performed according to vendor’s protocol. We preprocess MRI and FET-PET images using BraTS Toolkit [24] resulting in images resolution of $240\times 240\times 155$ with segmented tumor regions. Given our assumption that surrounding tissues remain static, we segment brain tissues based on an atlas registration [15].

### v4

@@ 2.1 Clinical Dataset Validation: Radiotherapy Planning and Recurrence Coverage
Our main analysis encompasses two distinct definitions of tumor recurrence: a broad definition including edema, enhancing core, and necrotic core, and a more specific definition aligned with current RANO guidelines [Chukwueke2019] focusing only on the enhancing core. Figures 5 and 6 present the mean and standard deviation of Recurrence Coverage for these definitions. Furthermore, we conduct a patient-specific comparison with the Standard Plan, classifying a model as ’Better’ if it provides higher coverage for an individual patient and ’Worse’ if it falls short of the Standard Plan’s coverage. In instances of equal coverage (e.g., both achieving 100%), we label the outcome as Equal. The comparison outcomes, including average results, are depicted in Figures 5c and 6c.
Noticeable disparities are evident between the outcomes from any segmentation recurrence and those from enhancing core recurrence only. First, let’s examine the results from radiotherapy targeting any segmentation recurrence. GliODIL demonstrates superior performance over uniform margin plans in most instances. Specifically, for the 58 patients with available FET-PET scans, the average Recurrence Coverage improved from 70.04% to 72.94%. For the 152 patients without FET-PET scans or where the scans were not incorporated into the model, the coverage increased from 63.59% to 67.80%. These improvements translate to a 35% and 38% difference in patient outcomes favoring GliODIL over the standard approach for the respective datasets.
Figure 5c reveals improvements, with the percentage of patients benefiting from the treatment increasing from 21% to 35% when comparing GliODIL, enhanced with FET-PET imaging, to the Standard Radiotherapy Plan with a 1.5cm margin. This emphasizes the value of integrating FET-PET modality into flexible models capable of utilizing multi-modal data. Conversely, incorporating this modality into models rigidly conforming to a basic predefined PDE family yields statistically insignificant changes in performance.
@@
In the analysis of enhancing core recurrence in follow-up scans, the hierarchy observed in the any segmentation recurrence study largely remains, with the exception that the Standard Plan demonstrates improved performance, outshining all models strictly based on PDE approaches. Instances of equal outcomes, as depicted in Figures 6c,d, become more prevalent, indicating that both compared methods often achieve 100% coverage.
GliODIL stands out as the only model consistently surpassing the Standard Plan. This leads to a marked increase in coverage improvement, from benefiting 5% of patients to 19%. The conclusion drawn is that it’s considerably more challenging for models to surpass the uniform margin around the pre-operative visible tumor. This difficulty arises because recurrences to the enhancing core typically occur closer to the resection cavity, rendering the migration paths, which these models leverage, less critical in this context compared to any segmentation recurrence. Any segmentation recurrences frequently occur in regions farther away, where anatomical features and topological barriers—factors not accounted for by a uniform margin—play a critical role. GliODIL, especially when incorporating FET-PET (see Figure 5), leverages additional information to refine predictions while maintaining a strong alignment with pre-operative data. As illustrated in the figure, this approach demonstrates improved recurrence coverage compared to the standard margin-based method.
Figure 5: Recurrence Coverage Analysis of Edema, Enhancing Core, and Necrotic Core in Real Patient Radiotherapy Plans a,b Recurrence coverage of selected volume radiotherapy plans. All radiotherapy plans have the same total volume. Output tumor cell distribution thresholds found through a grid search to match the Standard Plan volumes. c,d Average Recurrence Coverage and direct patient-by-patient comparisons to the Standard Plan.
Figure 6: Recurrence Coverage Analysis of the Enhancing Core in Real Patient Radiotherapy Plans. a,b Recurrence coverage of selected volume radiotherapy plans. All radiotherapy plans have the same total volume. Output tumor cell distribution thresholds found through a grid search to match the Standard Plan volumes. c,d Average Recurrence Coverage and direct patient-by-patient comparisons to the Standard Plan.
@@
Figure 7: Case-by-Case Comparison of Real Patient Data and GliODIL Limitations a Comparison of GliODIL with the Standard Plan across individual cases. b, c Examples of cases where GliODIL underperforms.
To support future advancements, we present a case-by-case comparison with the Standard Plan, highlighting instances where GliODIL performs suboptimally in Figure 7. In Figure 7b, incorporating brain fiber-related diffusion (as in [Buti_2024]) into the PDE constraint could potentially improve performance, as tumor recurrence appears to follow the orientation of brain fibers. In Figure 7c, the tumor’s multifocal characteristics exceed the local multifocal scope within which GliODIL was calibrated. A more sophisticated initial condition setup could potentially enhance GliODIL’s handling of such rare complex cases.
Ending this section, we note that the $\text{PDE}_{\text{GliODIL}}$ model, through the comparative analysis detailed in Section 2, aligns with the $\text{PDE}_{\text{CMA-ES}}$ in capturing pre-operative data yet surpasses it in recurrence prediction accuracy. This suggests the parameter estimations by $\text{PDE}_{\text{GliODIL}}$ more accurately reflect patient-specific conditions. Such improvements imply the data-driven component of GliODIL benefits the fine-tuning of PDE parameters which can be used for both diagnostics and forcasting.
@@ 3 Discussion
Leveraging its ability to synthesize information from diverse modalities, GliODIL significantly benefits from the integration of FET-PET imaging, which notably enhances its predictive accuracy beyond the capabilities of the Standard Radiotherapy Plan. This plan typically employs uniform safety margins around the visible tumor regions identified on MRI scans, aiming to delineate both the visible tumor and potential microscopic tumor extensions not visible on imaging. This approach, while clinically practiced, lacks the specificity required for optimal therapeutic outcomes. The improvements GliODIL brings to the table, as evidenced by the patient outcomes depicted in Figures 5 and 6, underscore the framework’s superior capability to utilize complex imaging data to refine treatment strategies. Such performance enhancements are not achieved by models strictly adhering to PDE solutions, which struggle to capture the full extent of tumor recurrence, even with the addition of extra imaging inputs.
GliODIL’s adaptive modeling approach provides a stark contrast to the Standard Radiotherapy Plan’s reliance on uniform margins, offering a much-needed precision in targeting the tumor. GliODIL has the potential to be extended by integrating additional diagnostic modalities beyond FET-PET imaging, such as MR diffusion, perfusion imaging, and MRSI metabolic imaging. These modalities could provide deeper insights into tissue properties like cellularity, microstructure, and metabolism, enhancing the model’s ability to represent complex tumor dynamics. Alternatively, using the same modalities, such as MRI and FET/PET, in a time-series format could provide significant benefits. Sequential imaging data would enable the calibration of more complex tumor growth models, which are otherwise limited by the ill-posed nature of single time-point data. This approach could offer a deeper understanding of tumor dynamics and improve model accuracy.
Our results thus highlight both the relevant contribution of advanced, biological imaging techniques to inform about the underlying tumor biology and the ability of our GliODIL framework to flexibly incorporate such additional information to improve inverse problem-solving validated by the clinically important tumor recurrence prediction task. Our current study is concentrated on immediate post-operative treatment, necessitating reliance on a single imaging time point.
Moreover, addressing uncertainties in parameter inference remains critical, and future work could explore embedding variational inference techniques to enhance the framework’s robustness.
@@ 4.2 Optimizing a Discrete Loss (ODIL)
The multi-grid ODIL technique, introduced in the paper [Karnakov2023], builds upon the original ODIL methodology by incorporating a multigrid decomposition scheme to fasten the convergence process. This technique is particularly adept at leveraging the multi-scale attributes of the forward problem. It decomposes the problem into various scale bands, each characterized by different levels of detail. Formally, for a uniform grid with dimensions $N_{1}=N$, a hierarchical sequence of coarser grids is introduced with sizes defined as $N_{i}=N/2^{i-1}$ for $i=1,\ldots,L$.
@@
As depicted in Figure 1,(3) which illustrates the multigrid domain, a tumor growth is simulated over an ensemble of Cartesian 4D grids in both time and space, with each grid level being coarser than the preceding one. This hierarchical decomposition allows the optimization algorithm to initially concentrate on the coarse-scale features, incrementally incorporating finer-scale details as the process evolves. Such an approach not only enables a more comprehensive exploration of the parameter space but also sidesteps the pitfalls of local minima and expedites convergence.
We focus the computational grids on the tumor region. This leads to an average resolution of $71\times 68\times 55$ for the area of interest in our images, significantly reduced from the original $240\times 240\times 155$. For parameter inference within the GliODIL framework, we employed a $48^{3}$ spatial resolution and 192 degrees in the temporal resolution. Our tests indicated that this resolution is sufficient, revealing no significant differences compared to inference using the native resolution. For a forward PDE run with inferred parameters by GliODIL, which we call $\text{PDE}_{\text{GliODIL}}$, we use the native resolution.
@@ 4.3 Imaging Model
$L_{\text{DATA}}=\lambda_{\text{CORE}}L_{\text{CORE}}(u,\theta_{\text{up}})+\lambda_{\text{EDEMA}}L_{\text{EDEMA}}(u,\theta_{\text{down}},\theta_{\text{up}})+\lambda_{\text{PET}}L_{\text{PET}}(u,s,\theta_{\text{BKG}})$ (10)
@@ 4.4 Final Loss Function
The final loss function measures discrepancy between proposed tumor cell concentrations $u(x,y,z,t)$ and our objective. It is a composite term that comprises contributions from multiple sources: the PDE constraint, data fitting, and additional regularization term. Specifically, the PDE loss, denoted by $L_{\text{PDE}}$ (introduced in 4.2), imposes the discretized PDE equation constraint. Initial condition loss term $L_{\text{IC}}$ in the overall loss function serves to enforce that the tumor at $t=0$ originates from a Gaussian origin (introduced in 4.1). The data loss (introduced in 4.3), denoted by $L_{\text{DATA}}=\lambda_{\text{CORE}}L_{\text{CORE}}+\lambda_{\text{EDEMA}}L_{\text{EDEMA}}+\lambda_{\text{PET}}L_{\text{PET}}$, accounts for matching tumor core and edema segmentations as well as fitting to PET metabolic data. Additional regularization term confines the inferred parameters within a plausible range $L_{\text{PARAMS}}$[Harpold2007].
@@
where $\lambda_{*}$ are detailed on in Section 2 experiments.
@@ 4.5 Evaluation Metrics
$\text{Dice Coefficient}=\frac{2\times|A\cap B|}{|A|+|B|}$ (18)
@@
The principal metric for evaluating our model’s effectiveness in radiotherapy planning is its accuracy in predicting tumor recurrence within the specified radiation volume after treatment. We identify the area of tumor recurrence using two definitions: a narrow definition focusing on the enhancing core observed in post-treatment patient data, and a broader definition that includes edema and the enhancing/necrotic core in the post-treatment follow-up MRI scans. We register the post-treatment patient anatomy to the pre-treatment to eliminate spatial shifts caused by surgeries. The Recurrence Coverage [%] metric refers to the percentage of segmented recurrences, that are encompassed within the radiation volume delineated by a given radiotherapy plan. Each radiotherapy plan maintains a uniform total volume, equivalent to that of the Standard Plan (CTV), which applies a consistent 1.5cm margin around the pre-operative segmented enhancing and necrotic regions. To facilitate a fair comparison, we adjust the Standard Plan’s total volume by excluding portions that extend beyond the patient’s anatomical boundaries before making any calculations.
4.6 Baselines
We compare our results with the Covariance Matrix Adaptation Evolution Strategy (CMA-ES) method [martin2021korali, weidner2024learnable], which relies on numerous simulations and employs a loss function from [lipkova2019personalized] to determine parameters that best fit the data. Its forward finite-difference scheme is analogous to that described in [jaroudi2020numerical]. In addition, we compare with the Learn-Morph-Infer (LMI) technique [ezhov2022learn], a deep learning framework for tumor growth model parameter estimation. We also include nnU-Net [isensee2018nnunet], the state-of-the-art neural network architecture for medical image segmentation, as an end-to-end data-driven baseline. For the standard nnU-Net experiments, the raw input features consisted of brain tissue distribution maps and pre-operative tumor segmentations. In contrast, for the nnU-Net-FET-PET variant, we additionally included FET-PET metabolic maps as input. The network outputs binary segmentation maps indicating predicted recurrence regions. To ensure compatibility with our radiotherapy planning pipeline and to standardize the output format with other models, we converted these binary outputs into a continuous representation using a Euclidean distance map. This transformation assigns a value of 1 to predicted recurrence regions, with values smoothly decreasing toward 0 as the distance from the predicted areas increases, resembling tumor cells distributionx. To maintain consistency between the GliODIL framework and nnU-Net results, we applied the same preprocessing steps across both methods. This included intensity normalization of input images, resampling to a common resolution, and brain mask application. The network was trained using an 80/20 train-test split. LMI employs data-driven convolutional neural networks to discover unknown PDE parameters, and CMA-ES is a state-of-the-art method [lipkova2019personalized] using both MRI and FET-PET data. We transitioned from TMCMC (Transitional Markov Chain Monte Carlo) to CMA-ES (Covariance Matrix Adaptation Evolution Strategy) to better scale our implementation for multiple patients. For clarity, solutions using the forward PDE finite difference method with parameters from CMA-ES and LMI are labeled as $\text{PDE}_{\text{CMA-ES}}$ and $\text{PDE}_{\text{LMI}}$, respectively, indicating direct PDE solutions.
4.7 Synthetic Dataset
@@
Clinical Dataset
[…]
To compare GliODIL-derived radiotherapy plans (clinical target volume, CTV) with current standard-of-care plans, we followed the ESTRO-EANO guidelines [Niyazi2023] to generate standard CTV maps. In brief, we dilated the preoperative tumor segmentation (tumor core and contrast-enhancing tumor) by a uniform margin of 15mm, excluding non-brain and CSF areas from the target volume.